By total count you determine that the concentration of an organism in a liquid culture is  ...

  1. By total count you determine that the concentration of an organism in a liquid culture is   5.6 x 107 CFU/ml and you have 1 ml of this culture. Assuming that total count equals viable count, design a practical dilution and plating scheme which will result in the formation of 280 colonies per plate. Include all steps and indicate volumes and dilutions.
Answer #1

CFU/ml = number of colonies / dilution factor *volume plated in ml.

Colony = 280 per plate.

Volume plated = 0.1ml

Dilution factor = Need to find out

5.6*107 = 280 /Dilution factor *0.1

Dilution factor = 280/ 5.6*107 *0.1

= 280/560*104

= 1/2*104

= 5/105

= 5*10-5

So the required dilution factor is 5*10-5

Please write to me if you need more help